# Why Policy Gradient Algorithms Work for Undiscounted Total-Reward MDPs

### v1

@@ 7 Experiments
For the experiments, we consider two toy examples: Frozenlake with $4\times 4$ states and 4 actions and CliffWalk with $3\times 7$ states and 4 actions. We use nonnegative rewards which ensures $V_{\mu}^{\star}=V_{+,\mu}^{\star}$ by Lemma 4, and uniform initial state distribution. Further details are provided in Appendix D.
@@
The results are shown in Figure 2. The natural policy gradient with adaptive step size exhibits the fastest convergence rate among the algorithms, as the guaranteed linear rate of Corollary 3 predicts. Note that both natural policy gradients converge to $V^{\star}_{\mu}$ while the projected policy gradient converges to $V_{\mu}^{\pi^{\star}_{\alpha}}$ for each $\alpha$, and smaller $\alpha$ makes projected policy gradient converge closer to $V_{\mu}^{\star}$ since $V_{\mu}^{\pi^{\star}_{\alpha}}$ increases monotonically to $V_{\mu}^{\star}$ as $\alpha\to 0$.
Additionally, we run an experiment with pathological MDP of Figure 1, shown in Appendix D.
@@ 8 Conclusion
In this work, we present the first analysis of policy gradient methods for undiscounted expected total-reward infinite-horizon MDPs. Our approach combines the classical recurrent-transient theory from Markov chain theory with prior analysis techniques for policy gradient methods. Specifically, we first establish invariance of the classification of MDP states on $\Pi_{+}$, where the value function is continuous, and define a new transient visitation measure that leads to a transient policy gradient. Based on this machinery, we establish non-asymptotic convergence rates for projected policy gradient and natural policy gradient in the undiscounted total-reward setting.
@@
Finally, we highlight that recurrent-transient classification of MDP states is a fundamental and broadly applicable technique. Previously, the recurrent-transient theory was also applied to improve the convergence of policy iteration independently of the discount factor (Fox & Landi, 1968; Bertsekas & Tsitsiklis, 1991; Scherrer, 2013). We expect that this technique can be used to analyze a wide range of RL algorithms in the undiscounted total-reward setting.
@@ Appendix A Omitted proofs in Section 4
For definitions of basic concepts of transient-recurrent theory such as irreducible class, communicating class, closedness, etc., please refer to Brémaud (2013, Chapter 2 and 3)
@@ A.1 Proof of Proposition 1
For any $\pi,\pi^{\prime}\in\Pi_{+}$, $P^{\pi}(s,s^{\prime})\neq 0$ if and only if $P^{\pi^{\prime}}(s,s^{\prime})\neq 0$ for $s,s^{\prime}\in\mathcal{S}$ by definition of $\Pi_{+}$. This implies $s$ and $s^{\prime}$ of $P^{\pi}$ communicate if and only if $s$ and $s^{\prime}$ of $P^{\pi^{\prime}}$ communicate, and thus communicating class is invariant for $\pi\in\Pi_{+}$. It was known that states in communicating class are all transient or recurrent (Brémaud, 2013, Theorem 3.1.6). Next, set of states is closed in $P^{\pi}$ if and only if it is closed in $P^{\pi^{\prime}}$. Therefore, since communicating class is closed if and only if it is recurrent in finite states MDP (Brémaud, 2013, Theorem 3.2.8), we obtain desired result. ∎
A.2 Proof of Lemma 1
@@
Since $\sum_{k=0}^{\infty}(P^{\pi})^{k}(s,s)=\infty$ for recurrent state $s$, $r^{\pi}(s)=0$ to satisfy Assumption 1. ∎
By Lemma 1 and 2, we directly obtain following Corollary.
Corollary 4.
Under Assumption 1, $V^{\pi}(s)=0$ for all recurrent states $s$.
A.3 Proof of Lemma 3
@@
By Proposition 1, we have fixed transient class and transient matrix $T^{\pi}$ is continuous on $\pi\in\Pi_{+}$. Since $(I-T^{\pi})^{-1}$ is also continuous with respect to $T^{\pi}$, by Lemma 2, $V^{\pi}$ and $V^{\pi}_{\mu}$ are continuous with respect to $\pi$. ∎
A.4 Proof of Lemma 5
@@
$V^{\pi}(s)-V^{\pi^{\prime}}(s)$ $=\sum_{a\in\mathcal{A}}Q^{\pi}(s,a)\pi(a\,|\,s)-\sum_{a\in\mathcal{A}}Q^{\pi^{\prime}}(s,a),\pi^{\prime}(a\,|\,s)$
$=\sum_{a\in\mathcal{A}}Q^{\pi^{\prime}}(s,\cdot)(\pi(\cdot\,|\,s)-\pi^{\prime}(\cdot\,|\,s))+\sum_{a\in\mathcal{A}}(Q^{\pi}(s,a)-Q^{\pi^{\prime}}(s,a))\pi(a\,|\,s)$
$=\sum_{a\in\mathcal{A}}Q^{\pi^{\prime}}(s,\cdot)(\pi(a\,|\,s)-\pi^{\prime}(a\,|\,s))+\sum_{s^{\prime}\in\mathcal{S}}\sum_{a\in\mathcal{A}}P(s^{\prime}\,|\,s,a)(V^{\pi}(s^{\prime})-V^{\pi^{\prime}}(s^{\prime})),\pi(a\,|\,s)$
@@
Corollary 5.
Under Assumption 1, if $\pi,\pi^{\prime}\in\Pi_{+}$,
$V_{\mu}^{\pi}-V_{\mu}^{\pi^{\prime}}$ $=\sum_{s^{\prime}\in\mathcal{S}}\sum_{a\in\mathcal{S}}Q^{\pi^{\prime}}(s^{\prime},a)(\pi(a\mid s^{\prime})-\pi^{\prime}(a\mid s^{\prime}))\delta^{\pi}_{\mu}(s^{\prime}).$
@@
By Lemma 1 and 2, we obtain $V^{\pi}(s)=0$ for any recurrent state $s$. Thus by Proposition 1, condition of Lemma 5 is satisfied. ∎
Corollary 6.
Under Assumption 1, if rewards are nonnegative, for $\pi\in\Pi$,
$V_{\mu}^{\star}-V_{\mu}^{\pi}$ $=\sum_{s^{\prime}\in\mathcal{S}}\sum_{a\in\mathcal{S}}Q^{\pi}(s^{\prime},a)(\pi^{\star}(a\mid s^{\prime})-\pi(a\mid s^{\prime}))\delta^{\pi^{\star}}_{\mu}(s^{\prime}).$
@@
Since $V^{\star}\geq V^{\pi}\geq 0$ by definition of $V^{\pi}$, $\pi$ satisfies condition of Lemma 5. ∎
A.5 Proof of Lemma 4
@@
A.6 Proof of Theorem 1
For the proof of Theorem 1, we first prove the following lemmas.
Lemma 8.
Under Assumption 1, for recurrent state $s$ of $P^{\pi}$ where $\pi\in\Pi_{+}$, $r(s,a)=0$.
@@ A.6 Proof of Theorem 1
If $r(s,a)\neq 0$ for some $a\in\mathcal{A}$, there exist $\pi\in\Pi_{+}$ such that $r^{\pi}(s)\neq 0$ since $\left\|{r}\right\|_{\infty}\leq R$. This is contradiction by Lemma 1. ∎
Now we prove Theorem 1.
@@
Fix $\pi\in\Pi_{+}$. We first clarify differentiability of $V^{\pi}$. For $\triangle\pi\in^{|\mathcal{S}|\times|\mathcal{A}|}$ which represents change of policy, we define $P^{\pi+\triangle\pi}(s,s^{\prime})=\sum_{a\in\mathcal{A}}(\pi+\triangle\pi)(s,a)P(s^{\prime}\,|\,s,a)$ and $r^{\pi+\triangle\pi}(s)=\sum_{a\in\mathcal{A}}(\pi+\triangle\pi)(s,a)r(s,a)$ for all $s,s^{\prime}\in\mathcal{S}$. Then, if $s$ is recurrent, $r^{\pi+\triangle\pi}(s)=0$ by Lemma 8.
By definition of $\Pi_{+}$, there exsit open ball $B(\pi,\epsilon)$ such that $(\pi+\triangle\pi)(a\,|\,s)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$ and $\triangle\pi\in B(\pi,\epsilon)$. Then, $P^{\pi}(s,s^{\prime})\neq 0$ if and only if $P^{\pi+\triangle\pi}(s,s^{\prime})\neq 0$ for all $s,s^{\prime}\in\mathcal{S}$, and we can define perturbed transient matrix $T^{\pi+\triangle\pi}$ for $\pi\in\Pi_{+}$.
Since spectral radius is continuous with respect to element of matrix (Horn & Johnson, 2012, Theorem 2.4.9.2), and $T^{\pi}$ is continuous with respect to $\pi\in\Pi_{+}$, there also exist open ball $B^{\prime}(\pi,\epsilon)\subset B(\pi,\epsilon)$ such that spectral radius of $T^{\pi+\triangle\pi}$ is smaller than $1$. (Note that this argument is valid since set of transient class and recurrent class fixed by Proposition 1.) Thus
@@
and this implies well-definedness of value function on $\pi+\triangle\pi$ where $c\in\Pi_{+}$ and $\triangle\pi\in B^{\prime}(\pi,\epsilon)$. Then, by Lemma 2, differentiability of $T^{\pi}$ and $r^{\pi}$ on $\pi\in\Pi_{+}$ implies differentiability of $V^{\pi}$, and it can be easily seen that $T^{\pi}$ is differentiable with respect to $\pi$ since each element of $P^{\pi}$ is differentiable and transient class is fixed by Proposition 1. $r^{\pi}$ is obviously differentiable.
@@
where third equality comes from matrix calculus $\frac{\partial A^{-1}}{\partial\theta}=A^{-1}\frac{\partial A}{\partial\theta}A^{-1}$ for $A(\theta)\in^{|\mathcal{S}|\times|\mathcal{S}|}$ and forth equality is from fact that $\Theta\in^{|\mathcal{S}|\times|\mathcal{S}||\mathcal{A}|}$ is matrix form of policy $\pi_{\theta}$ satisfying $\Theta P=P^{\pi_{\theta}}$ and $\Theta r=r^{\pi_{\theta}}$ and $\frac{\partial T^{\pi_{\theta}}}{\partial\theta}V^{\pi_{\theta}}=\frac{\partial P^{\pi_{\theta}}}{\partial\theta}V^{\pi_{\theta}}$ by Corollary 4 and Proposition 1. ∎
@@ B.1 Proof of Lemma 6
We basically follow the proof strategy of Agarwal et al. (2021); Mei et al. (2020) . Let $\theta_{\beta}=\theta+\beta u$. Then, with direct parametrization,
@@
We now prove another key lemma for Theorem 2.
Lemma 9 (Gradient domination).
Under Assumption 1, for $\pi\in\Pi_{\alpha}$,
@@
We basically follow the proof strategy of Agarwal et al. (2021). Let $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)$. Then, we have
@@
where first equality comes from Corollary 5, the last inequality follows since $\max_{\bar{\pi}\in\Pi_{\alpha}}\left(\sum_{a}\bar{\pi}(a\,|\,s)A^{\pi}(s,a)\right)\geq 0$ for all $s\in\mathcal{S}$ and policies $\pi\in\Pi_{\alpha}$. Also, we have
@@
where the first equality follows from the fact that $\max_{\bar{\pi}}$ is attained at an action which maximizes $A^{\pi}(s,\cdot)$, the second equality is from $\sum_{a}\pi(a|s)A^{\pi}(s,a)=0$, the third equality is from $\sum_{a}(\bar{\pi}(a|s)-\pi(a|s))V^{\pi}(s)=0$ for all $s$, and the last equality follows from the Theorem 1 with direct parameterization. Finally,
@@
B.2 Proof of Theorem 2
Following the proof strategy of Xiao (2022), we consider composite optimization problem:
@@ B.2 Proof of Theorem 2
where $f$ is $L$-smooth and $\Psi$ is is proper, convex, and closed (Rockafellar, 1970). Define $F^{\star}=\min_{x}F(x)$.
@@
Specifically, let $\eta_{k}=\frac{1}{L}.$ and define
@@
Definition (weak gradient-mapping domination).
@@
(Nesterov, 2013, Theorem 1) Consider the composite optimization problem where $f$ is $L$-smooth on closed convex set $C$ and $\Psi$ is indicator function with $C$. Then, for $x,y\in C$,
@@
Note that if $\Psi$ is indicator function with convex closed non empty subset of $C\in^{n}$, $\Psi$ is closed, convex, and proper and $\textbf{prox}_{\eta\Psi}=\textbf{proj}_{C}$ where proj is projection operator.
@@
To prove Theorem 2, we first prove following lemma.
Lemma 10.
Under Assumption 1, for a given $\mu$ with full support, suppose that $V^{\pi}_{\mu}$ is $L$-smooth for $\pi\in\Pi_{\alpha}$. Then,
@@
Combining the above bound with Lemma 9 gives the desired inequality. ∎
Now we are ready to Theorem 2.
@@
By lemma 10, weak gradient-mapping domination condition holds with
@@ B.3 Proof of Corollary 1
Let $\alpha=\frac{\epsilon}{2|\mathcal{S}||\mathcal{A}|\left\|{\delta^{\pi^{\star}}_{\mu}}\right\|_{\infty}\left\|{Q^{\pi^{\star}}}\right\|_{\infty}}$. Since there exist $\pi\in\Pi_{\alpha}$ such that $\left\|{\pi^{\star}-\pi}\right\|_{\infty}\leq\alpha$, by Corollary 6, we have
$V_{\mu}^{\star}-V_{\mu}^{\pi}$ $=\sum_{s^{\prime}\in\mathcal{S}}\sum_{a\in\mathcal{S}}Q^{\pi}(s^{\prime},a)(\pi^{\star}(a\mid s^{\prime})-\pi(a\mid s^{\prime}))\delta^{\pi^{\star}}_{\mu}(s^{\prime})$
@@
and this implies $V^{\star}_{\mu}\;-\;V^{\pi^{\star}_{\alpha}}_{\mu}\leq\frac{\epsilon}{2}.$ and the result comes from Theorem 2 by having $V^{\pi^{\star}_{\alpha}}_{\mu}-V^{\pi_{k}}_{\mu}\leq\frac{\epsilon}{2}$. ∎

### v2

@@ 8 Experiments
For the experiments, we consider two toy examples: Frozenlake with $4\times 4$ states and 4 actions and CliffWalk with $3\times 7$ states and 4 actions. We use nonnegative rewards which ensures $V_{\mu}^{\star}=V_{+,\mu}^{\star}$ by Lemma 4.5, and uniform initial state distribution. Further details are provided in Appendix E.
@@
The results are shown in Figure 2. The natural policy gradient with adaptive step size exhibits the fastest convergence rate among the algorithms, as the guaranteed linear rate of Corollary 6.5 predicts. Note that both natural policy gradients converge to $V^{\star}_{\mu}$ while the projected policy gradient converges to $V_{\mu}^{\pi^{\star}_{\alpha}}$ for each $\alpha$, and smaller $\alpha$ makes projected policy gradient converge closer to $V_{\mu}^{\star}$ since $V_{\mu}^{\pi^{\star}_{\alpha}}$ increases monotonically to $V_{\mu}^{\star}$ as $\alpha\to 0$.
Additionally, we run an experiment with pathological MDP of Figure 1, shown in Appendix E.
@@ 9 Conclusion
In this work, we present the first analysis of policy gradient methods for undiscounted expected total-reward infinite-horizon MDPs. Our approach combines the classical recurrent-transient theory from Markov chain theory with prior analysis techniques for policy gradient methods. Specifically, we first establish invariance of the classification of MDP states on $\Pi_{+}$, where the value function is continuous, and define a new transient visitation measure that leads to a transient policy gradient. Based on this machinery, we establish non-asymptotic convergence rates for projected policy gradient and natural policy gradient in the undiscounted total-reward setting. Lastly, we present the sample complexity of stochastic natural policy gradient by approximating the $Q$-function with a generative model.
@@
Finally, we highlight that recurrent-transient classification of MDP states is a fundamental and broadly applicable technique. Previously, the recurrent-transient theory was also applied to improve the convergence of policy iteration independently of the discount factor (Fox and Landi, 1968; Bertsekas and Tsitsiklis, 1991; Scherrer, 2013). We expect that this technique can be used to analyze a wide range of RL algorithms in the undiscounted total-reward setting.
@@ Appendix A Omitted proofs in Section 4
For definitions of basic concepts of transient-recurrent theory such as irreducible class, communicating class, closedness, etc., please refer to Brémaud (2013, Chapters 2 and 3)
@@ A.1 Proof of Proposition 1
For any $\pi,\pi^{\prime}\in\Pi_{+}$, $P^{\pi}(s,s^{\prime})\neq 0$ if and only if $P^{\pi^{\prime}}(s,s^{\prime})\neq 0$ for $s,s^{\prime}\in\mathcal{S}$ by definition of $\Pi_{+}$. This implies $s$ and $s^{\prime}$ of $P^{\pi}$ communicate if and only if $s$ and $s^{\prime}$ of $P^{\pi^{\prime}}$ communicate, and thus communicating class is invariant for $\pi\in\Pi_{+}$. It is known that states in communicating class are all transient or recurrent (Brémaud, 2013, Theorem 3.1.6). Next, the set of states is closed in $P^{\pi}$ if and only if it is closed in $P^{\pi^{\prime}}$. Therefore, since communicating class is closed if and only if it is recurrent in finite states MDP (Brémaud, 2013, Theorem 3.2.8), we obtain the desired result. ∎
A.2 Proof of Lemma 4.2
@@
Since $\sum_{k=0}^{\infty}(P^{\pi})^{k}(s,s)=\infty$ for recurrent state $s$, $r^{\pi}(s)=0$ to satisfy Assumption 2.1. ∎
By Lemma 4.2 and 4.3, we directly obtain the following Corollary.
Corollary A.1.
Under Assumption 2.1, $V^{\pi}(s)=0$ for all recurrent states $s$.
A.3 Proof of Lemma 4.4
@@
$\pi\mapsto P^{\pi}$ is continuous, and by Proposition 1, $\pi\mapsto T^{\pi}$ is also continuous on $\Pi_{+}$. Since $(I-T^{\pi})^{-1}$ is continuous with respect to $T^{\pi}$, by Lemma 4.3, $V^{\pi}$ and $V^{\pi}_{\mu}$ are continuous with respect to $\pi$. ∎
A.4 Proof of Lemma 4.6
@@
$V^{\pi}(s)-V^{\pi^{\prime}}(s)$ $=\sum_{a\in\mathcal{A}}Q^{\pi}(s,a)\pi(a\,|\,s)-\sum_{a\in\mathcal{A}}Q^{\pi^{\prime}}(s,a)\pi^{\prime}(a\,|\,s)$
$=\sum_{a\in\mathcal{A}}Q^{\pi^{\prime}}(s,a)(\pi(a\,|\,s)-\pi^{\prime}(a\,|\,s))+\sum_{a\in\mathcal{A}}(Q^{\pi}(s,a)-Q^{\pi^{\prime}}(s,a))\pi(a\,|\,s)$
$=\sum_{a\in\mathcal{A}}Q^{\pi^{\prime}}(s,a)(\pi(a\,|\,s)-\pi^{\prime}(a\,|\,s))+\sum_{s^{\prime}\in\mathcal{S}}\sum_{a\in\mathcal{A}}P(s^{\prime}\,|\,s,a)(V^{\pi}(s^{\prime})-V^{\pi^{\prime}}(s^{\prime}))\pi(a\,|\,s)$
@@
Corollary A.2.
Under Assumption 2.1, if $\pi,\pi^{\prime}\in\Pi_{+}$,
$V_{\mu}^{\pi}-V_{\mu}^{\pi^{\prime}}$ $=\sum_{s^{\prime}\in\mathcal{S}}\sum_{a\in\mathcal{A}}Q^{\pi^{\prime}}(s^{\prime},a)(\pi(a\mid s^{\prime})-\pi^{\prime}(a\mid s^{\prime}))\delta^{\pi}_{\mu}(s^{\prime}).$
@@
By Lemma 4.2 and 4.3, we obtain $V^{\pi}(s)=0$ for any recurrent state $s$. Thus by Proposition 1, condition of Lemma 4.6 is satisfied. ∎
Corollary A.3.
Under Assumption 2.1, if rewards are nonnegative, for $\pi\in\Pi$,
$V_{\mu}^{\star}-V_{\mu}^{\pi}$ $=\sum_{s^{\prime}\in\mathcal{S}}\sum_{a\in\mathcal{A}}Q^{\pi}(s^{\prime},a)(\pi^{\star}(a\mid s^{\prime})-\pi(a\mid s^{\prime}))\delta^{\pi^{\star}}_{\mu}(s^{\prime}).$
@@
Since $V^{\star}\geq V^{\pi}\geq 0$ by definition of $V^{\pi}$, $\pi$ satisfies condition of Lemma 4.6. ∎
A.5 Proof of Lemma 4.5
@@
A.6 Proof of Theorem 4.7
For the proof of Theorem 4.7, we first prove the following lemmas.
Lemma A.4.
Under Assumption 2.1, for recurrent state $s$ of $P^{\pi}$ where $\pi\in\Pi_{+}$, $r(s,a)=0$.
@@
If $r(s,a)\neq 0$ for some $a\in\mathcal{A}$, there exists $\pi\in\Pi_{+}$ such that $r^{\pi}(s)\neq 0$ since $\left\|{r}\right\|_{\infty}\leq R$. This is a contradiction by Lemma 4.2. ∎
Now we prove Theorem 4.7.
@@
Fix $\pi\in\Pi_{+}$. We first clarify differentiability of $V^{\pi}$. For $\triangle\pi\in^{|\mathcal{S}|\times|\mathcal{A}|}$ which represents change of policy, we define $P^{\pi+\triangle\pi}(s,s^{\prime})=\sum_{a\in\mathcal{A}}(\pi+\triangle\pi)(s,a)P(s^{\prime}\,|\,s,a)$ and $r^{\pi+\triangle\pi}(s)=\sum_{a\in\mathcal{A}}(\pi+\triangle\pi)(s,a)r(s,a)$ for all $s,s^{\prime}\in\mathcal{S}$. Then, if $s$ is recurrent, $r^{\pi+\triangle\pi}(s)=0$ by Lemma A.4.
By definition of $\Pi_{+}$, there exist an open ball $B(\pi,\epsilon)$ such that $(\pi+\triangle\pi)(a\,|\,s)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$ and $\triangle\pi\in B(\pi,\epsilon)$. Then, $P^{\pi}(s,s^{\prime})\neq 0$ if and only if $P^{\pi+\triangle\pi}(s,s^{\prime})\neq 0$ for all $s,s^{\prime}\in\mathcal{S}$, and we can define perturbed transient matrix $T^{\pi+\triangle\pi}$ for $\pi\in\Pi_{+}$.
Since spectral radius is continuous with respect to entries of matrix (Horn and Johnson, 2012, Theorem 2.4.9.2), and $T^{\pi}$ is continuous with respect to $\pi\in\Pi_{+}$, there also exist open ball $B^{\prime}(\pi,\epsilon)\subset B(\pi,\epsilon)$ such that spectral radius of $T^{\pi+\triangle\pi}$ is smaller than $1$. (Note that this argument is valid since set of transient class and recurrent class fixed by Proposition 1.) Thus
@@
and this implies well-definedness of value function on $\pi+\triangle\pi$ where $c\in\Pi_{+}$ and $\triangle\pi\in B^{\prime}(\pi,\epsilon)$. Then, by Lemma 4.3, differentiability of $T^{\pi}$ and $r^{\pi}$ on $\pi\in\Pi_{+}$ implies differentiability of $V^{\pi}$, and it can be easily seen that $T^{\pi}$ is differentiable with respect to $\pi$ since each element of $P^{\pi}$ is differentiable and transient class is fixed by Proposition 1. $r^{\pi}$ is obviously differentiable.
@@
where third equality comes from matrix calculus $\frac{\partial A^{-1}}{\partial\theta}=A^{-1}\frac{\partial A}{\partial\theta}A^{-1}$ for $A(\theta)\in^{|\mathcal{S}|\times|\mathcal{S}|}$ and forth equality is from fact that $\Theta\in^{|\mathcal{S}|\times|\mathcal{S}||\mathcal{A}|}$ is matrix form of policy $\pi_{\theta}$ satisfying $\Theta P=P^{\pi_{\theta}}$ and $\Theta r=r^{\pi_{\theta}}$ and $\frac{\partial T^{\pi_{\theta}}}{\partial\theta}V^{\pi_{\theta}}=\frac{\partial P^{\pi_{\theta}}}{\partial\theta}V^{\pi_{\theta}}$ by Corollary A.1 and Proposition 1. ∎
@@ Appendix B Omitted proofs in Section 5
We basically follow the proof strategy of (Agarwal et al., 2021; Mei et al., 2020) . Let $\theta_{\beta}=\theta+\beta u$. Then, with direct parametrization,
@@
We now prove another key lemma for Theorem 5.2.
Lemma B.1 (Gradient domination).
Under Assumption 2.1, for $\pi\in\Pi_{\alpha}$,
@@
We basically follow the proof strategy of (Agarwal et al., 2021). Let $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)$. Then, we have
@@
where first equality comes from Corollary A.2, the last inequality follows since $\max_{\bar{\pi}\in\Pi_{\alpha}}\left(\sum_{a}\bar{\pi}(a\,|\,s)A^{\pi}(s,a)\right)\geq 0$ for all $s\in\mathcal{S}$ and policies $\pi\in\Pi_{\alpha}$. Also, we have
@@
where the first equality follows from the fact that $\max_{\bar{\pi}}$ is attained at an action which maximizes $A^{\pi}(s,\cdot)$, the second equality is from $\sum_{a}\pi(a|s)A^{\pi}(s,a)=0$, the third equality is from $\sum_{a}(\bar{\pi}(a|s)-\pi(a|s))V^{\pi}(s)=0$ for all $s$, and the last equality follows from the Theorem 4.7 with direct parameterization. Finally,
@@
B.2 Proof of Theorem 5.2
Following the proof strategy of (Xiao, 2022), we consider composite optimization problem:
@@
where $f$ is $L$-smooth and $\Psi$ is proper, convex, and closed (Rockafellar, 1970). Define $F^{\star}=\min_{x}F(x)$.
@@
Specifically, let $\eta_{k}=\frac{1}{L}$, and define
@@
Definition B.2 (weak gradient-mapping domination).
@@
(Nesterov, 2013, Theorem 1) Consider the composite optimization problem where $f$ is $L$-smooth on closed convex set $C$ and $\Psi$ is the indicator function with $C$. Then, for $x,y\in C$,
@@
Note that if $\Psi$ is the indicator function of a convex closed non empty subset of $C\in^{n}$, $\Psi$ is closed, convex, and proper and $\textbf{prox}_{\eta\Psi}=\textbf{proj}_{C}$ where proj is projection operator.
@@
To prove Theorem 5.2, we first prove following lemma.
Lemma B.3.
Under Assumption 2.1, for a given $\mu$ with full support, suppose that $V^{\pi}_{\mu}$ is $L$-smooth for $\pi\in\Pi_{\alpha}$. Then,
@@
Combining the above bound with Lemma B.1 gives the desired inequality. ∎
Now we are ready to prove Theorem 5.2.
@@
By lemma B.3, weak gradient-mapping domination condition holds with
@@
Let $\alpha=\frac{\epsilon}{2|\mathcal{S}||\mathcal{A}|\left\|{\delta^{\pi^{\star}}_{\mu}}\right\|_{\infty}\left\|{Q^{\pi^{\star}}}\right\|_{\infty}}$. Since there exist $\pi\in\Pi_{\alpha}$ such that $\left\|{\pi^{\star}-\pi}\right\|_{\infty}\leq\alpha$, by Corollary A.3, we have
$V_{\mu}^{\star}-V_{\mu}^{\pi}$ $=\sum_{s^{\prime}\in\mathcal{S}}\sum_{a\in\mathcal{A}}Q^{\pi}(s^{\prime},a)(\pi^{\star}(a\mid s^{\prime})-\pi(a\mid s^{\prime}))\delta^{\pi^{\star}}_{\mu}(s^{\prime})$
@@
and this implies $V^{\star}_{\mu}\;-\;V^{\pi^{\star}_{\alpha}}_{\mu}\leq\frac{\epsilon}{2}.$ and the result comes from Theorem 5.2 by having $V^{\pi^{\star}_{\alpha}}_{\mu}-V^{\pi_{k}}_{\mu}\leq\frac{\epsilon}{2}$. ∎